# Alleviating Non-identifiability: a High-fidelity Calibration Objective for Financial Market Simulation with Multivariate Time Series Data

### v2

@@ 3.2 Alleviating the Non-identifiability with Multiple Features
Note that, given any two parameters $\omega_{1},\omega_{2}\in\Omega^{D,k}$, only they are simultaneously non-identifiable in all $K$ individual calibration tasks, i.e., $\omega_{1},\omega_{2}\in\{\mathcal{S}^{D,k}\}_{k=1}^{K}$, are they non-identifiable with the general calibration task with respect to $\hat{\mathbf{X}}_{T}$. Otherwise, their discrepancy to $\hat{\mathbf{X}}_{T}$ can be distinguishable in at least one feature and thus are identifiable. That is to say, only the intersection of $\{\mathcal{S}^{D,k}\}_{k=1}^{K}$ can be regarded as the non-identifiable set in the multivariate time series setting. Hence, the non-identifiability in the multivariate time series setting can be calculated as:
$P(\omega\in\{\mathcal{S}^{D,k}\}_{k=1}^{K}|\hat{\mathbf{X}}_{T})=\frac{|\cap_{k=1}^{K}\mathcal{S}^{D,k}|}{|\Omega^{D}|}$ (3)
It is easy to obtain that $\frac{|\cap_{k=1}^{K}\mathcal{S}^{D,k}|}{|\Omega^{D}|}\leq\frac{|\mathcal{S}^{D,k}|}{|\Omega^{D,k}|},k=1,..,K$. Hence, we have:
$P(\omega\in\{\mathcal{S}^{D,k}\}_{k=1}^{K}|\hat{\mathbf{X}}_{T})\leq P(\omega\in\mathcal{S}^{D,k}|\hat{\mathbf{X}}_{T}^{k})$ (4)
In Eq.(3) and Eq.(4), the equality holds only if the $K$ non-identifiable sets are fully overlapped. This can be caused by two reasons. First, the $K$ features are fully dependent that the time series data of one feature can be deduced from other features. This suggests that the selected features should be as diverse as possible. Second, the parameter spaces of all the other $K$ individual tasks are full of non-identifiable sets, i.e., $\mathcal{S}^{D,k}=\Omega^{D,k},k=1,2,...,K$, which is highly unlikely with existing discrepancy metrics $D$. Therefore, with proper choice of the features and $D$, the non-identifiability of calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$ is smaller than that of the univariate $\hat{\mathbf{X}}_{T}^{k}$. In other words, calibrating multiple features can alleviate the non-identifiability issue. Next, we estimate how fast it decreases with $K$ as follows.
For clarity, let $A_{k}$ denote the event of a uniformly randomly sampled parameter falling into the non-identifiable set of the $k$-th individual calibration task. The probability of happening $A_{k}$ is $P(A_{k})=P(\omega\in\mathcal{S}^{D,k}|\hat{\mathbf{X}}_{T}^{k})$. Then the non-identifiability in the multivariate time series setting can be rewritten as:
$P(\omega\in\{\mathcal{S}^{D,k}\}_{k=1}^{K}|\hat{\mathbf{X}}_{T})$
$=$ $P(A_{1},A_{2},...,A_{K})$
$=$ $P(A_{1})P(A_{2}|A_{1})P(A_{3}|A_{2},A_{1})\cdots P(A_{K}|A_{K-1},...,A_{1})$
@@ 3.3 Aggregating Multiple Features via Maximization
Note that, the above discussions require the multivariate time series $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$ being calibrated separately as $K$ univariate time series with existing discrepancy metrics $D$. This results in $K$ discrepancy values of $D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),k=1,\cdots,K$. This section discusses how to optimally aggregate the $K$ values, so that the practical alleviation of the non-identifiability follows the above analysis.
Let $F$ denote the proposed objective function for calibrating the multivariate $\hat{\mathbf{X}}_{T}$. To achieve $\mathcal{S}^{F}=\cap_{k=1}^{K}\mathcal{S}^{D,k}$, the proposed way is to ensure every location in $\Omega^{F}$ to be the worst among all $K$ parameter space $\{\Omega^{D,k}\}_{k=1}^{K}$ (see Figure  3 for illustration). In this case, any parameter $\omega$ that is not simultaneously in all $K$ non-identifiable sets will make it fall outside of the $\mathcal{S}^{F}$. This is because that the parameters outside the non-identifiable set is worst than those who are non-identifiable from the ground-truth parameter $\omega^{*}$.
Given that the calibration problem is a minimization problem, the new objective function $F$ can be implemented with a maximization aggregation among $K$ individual calibration tasks with $D$. That is,
$\min_{\omega\in\Omega}F(\hat{\mathbf{X}}_{T},M(w)=\mathbf{X}_{T})$ (6)
$=\min_{\omega\in\Omega}$ $\max_{k\in\{1,\cdots,K\}}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k})$
Figure 3: Optimizing the maximization of two functions is equivalent to the search in the intersection between the two search spaces.
@@ 4 Empirical Studies
This section empirically verifies two major claims of this paper.
  • With the increasing number of features applied in the calibration, the non-identifiable set shrinks largely.
@@ 4.1 Experimental Settings
$D(\hat{\mathbf{X}}_{T}^{k},\mathbf{X}_{T}^{k})=\int_{-\infty}^{+\infty}|U(x)-V(x)|\rm{d}\textit{x}$ (7)
where
$U(x)=\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\hat{\mathbf{X}}_{t}^{k}\leq x)$ (8)
$V(x)=\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\mathbf{X}_{t}^{k}\leq x)$
Here $U$ and $V$ are the empirical cumulative distribution functions for the data set $\hat{\mathbf{X}}_{T}^{k}$ and $\mathbf{X}_{T}^{k}$, respectively. And $\mathcal{I}(\cdot)$ is an indicator function that returns 1 if the input event is true, otherwise it returns 0.
The simulation model. This work considers the Preis-Golke-Paul-Schneid (PGPS) model as the simulation model, due to its popularity in recent financial market simulation works  [11, 6]. The PGPS model simulates the market dynamics through interactions between two types of agents, i.e., 125 liquidity providers and 125 liquidity takers. At the $t$-th time-step, each liquidity provider submits a limited order at a fixed probability $\alpha$ with the default volume equals 1. The probability of a limited order being bid side or ask side equals to 0.5. Each liquidity taker submits a market order at a fixed probability $\mu$ with the default volume equals 1. The probability of a market order being bid side or ask side is $q_{taker}(t)$ or $1-q_{taker}(t)$, respectively. The probability $q_{taker}(t)$ is specified by a mean-reverting random walk with mean equals 0.5, mean reversion probability equals $0.5+|q_{taker}(t)-0.5|$, and the increment sizes towards mean equals ±$\pm\Delta_{s}$. Moreover, the liquidity taker has a probability $\delta$ to cancel its untraded limited order. Let $p_{a}(t)$ and $p_{b}(t)$ represents the best ask and bid price, respectively. The price of a market order is determined by the market, i.e., the best price at the opposite side of the LOB. The price of a limited order is determined as $p_{a}(t)+(\lambda(t)\log(u))+s$, where $s=1$ if it is a ask order and otherwise $s=-1$, and $u\sim U(0,1)$. Here $\lambda(t)$ is a time-variant order placement depth parameter and is calculated as
$\lambda(t)=\lambda_{0}(1+\frac{|q_{taker}(t)-\frac{1}{2}|}{\sqrt{<q_{taker}(t)-\frac{1}{2}>^{2}}}C_{\lambda})$ (9)
$\sqrt{<q_{taker}(t)-\frac{1}{2}>^{2}}$ is a pre-computed value. Before initiating the simulation, it is computed in advance through $10^{5}$ Monte Carlo iterations, ensuring its convergence to the accurate standard deviation.
To summarize, the PGPS model has 6 hyper-parameters shared by all agents that need to be calibrated, i.e., $\omega=[\delta,\lambda_{0},C_{\lambda},\Delta_{s},\alpha,\mu]$.
The optimization method. Since the order matching process needs to align with the real market, the PGPS model should be run on the executable programs based some trading simulators [^2] This work builds the simulation environment based on the Multi-Agent Exchange Environment (MAXE) simulator developed by University of Oxford  [27].. As a result, the calibration objective function is non-differentiable and the calibration process often has to resort to non-derivative optimization methods. In such contexts, both Bayesian optimization methods and evolutionary algorithms are particularly suitable. This work employs the well-known Particle Swarm Optimization (PSO) algorithm  [28] to optimize $\omega$ with respect to Eq.(6). For simplicity, the standard version of PSO  [29] is considered. Without further fine-tuning, the hyper-parameters of PSO follow the suggested configurations, where the population size is set to 40, the inertia weight is set to 0.8, the cognitive and social crossover parameters $c_{1}=0.5,c_{2}=0.5$.
@@
  • $f_{3}$, the mid-price return within each $t$-th step $\ln m_{t+1}-\ln m_{t}$;
@@
Parameters $\mathbf{\delta}$ $\mathbf{\lambda_{0}}$ $\mathbf{C_{\lambda}}$ $\mathbf{\Delta_{S}}$ $\mathbf{\alpha}$ $\mathbf{\mu}$
data 1 0.025 100 10 0.001 0.15 0.025
data 2 0.002 200 10 0.002 0.1 0.03
data 3 0.05 152 45 0.003 0.13 0.05
data 4 0.02 288 27 0.003 0.11 0.04
data 5 0.04 62 35 0.003 0.12 0.04
data 6 0.01 171 15 0.0015 0.15 0.03
data 7 0.033 114 14 0.0033 0.1 0.047
data 8 0.01 129 30 0.0017 0.05 0.03
data 9 0.02 62 2 0.003 0.14 0.02
data 10 0.01 87 23 0.001 0.09 0.05
Table 1: Parameter settings of generating 10 synthetic data with the PGPS model.
The test protocol. The test data contains 11 targeted time series to be calibrated, including 10 synthetic data and 1 real data. The 10 synthetic time series are generated by setting 10 different parameters of the PGPS model as shown in Table  1, each of which consists of 3600 time steps at the frequency of 1 second. The real data is chosen as 000001.sz from Shenzhen Stock Exchange of China, consisting of 1200 time steps at a frequency of 3 seconds from 9:30 a.m. to 10:30 a.m. of a day in 2019. The time budget of $10^{4}$ simulation evaluations is allowed for the calibration of each targeted data. That is, the PSO runs for 250 iterations with population size 40, and the best parameter with the lowest objective function value is applied into PGPS model to simulate data as $\mathbf{X}_{T}$. The performance of the calibration is quantitatively validated using the Wasserstein (W) distance and the Mean Square Error (MSE) commonly seen in time series analysis. Both metrics are applied to measure the discrepancy between the synthetic data and simulated data on each feature, normalized with the min-max method, respectively.
(a) $F_{1}$
(b) $F_{2}$
(c) $F_{3}$
(d) $F_{4}$
(e) $F_{5}$
(f) $F_{6}$
Figure 4: With the increasing number of features, the cardinality of the non-identifiable set decreases significantly.
@@ 4.2 Results and Analysis
Effectiveness of multiple features. To show how the increasing numbers of features alleviate the non-identifiability issue, we depict the changes of the 2-dimensional parameter space of $[\alpha,\mu]$. Specifically, a target data of 3600 time steps is generated using the PGPS model with the commonly recommended default parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$  [6]. On this default setting, 10000 combinations are created by sampling in $[\alpha,\mu]$ in grid. By calculating the objective value $F$ between the target data and each of the 10000 combinations, it obtains the objective values of the 10000 combinations in the space of $[\alpha,\mu]$, which are depicted as Figure  4.(a)-(f). The 6 figures in Figure  4.(a)-(f) are depicted according to the objective values of $F_{1}$, $F_{2}$,…,$F_{6}$, respectively. Only the combinations whose objective values rank in the best 10% over all the corresponding features, are they depicted in blue, otherwise they will be depicted in yellow. For example, in Figure  4.(c), the combinations depicted in blue are those whose discrepancy values rank in 10% among the 10000 combinations on all the first 3 features $\{f_{i}\}_{i=1}^{3}$. By this means, the blue points indicate the intersection of the combinations who rank best 10% in all the corresponding features in $F$. In this regard, the non-identifiable set are also depicted in blue as they are indistinguishable with $\omega^{*}$.
As can be seen that, with the increaseha of the features, the blue set, as well as the non-identifiable set, decreases significantly. When 5 to 6 features are included in the objective function, there leaves very few blue points, indicating that the non-identifiability issue is indeed alleviated very effectively.
@@
Note that, although the non-identifiable set of any $F_{i}$ is verified to be smaller than that of any $F_{j}$, $i>j$, it does not deterministically lead to better calibration performance of $F_{i}$ than $F_{j}$. Because $F$ mainly decides the parameter space, while the calibration performance is also jointly influenced by the optimization algorithms. Given the calibration problem is black-box with multiple local optima, it is non-trivial for the optimization algorithms to find the global optimum even the parameter space is modeled fully identifiable.
data $obj$ $F_{1}$ $F_{2}$ $F_{3}$ $F_{4}$ $F_{5}$ $F_{6}$
Data 1 W 7.49 3.17 4.85 6.25 2.59 4.90
MSE 6.97 4.46 3.93 6.25 4.38 4.04
Data 2 W 3.68 1.58 2.60 5.81 4.40 3.88
MSE 4.15 4.18 4.88 5.25 5.15 4.33
Data 3 W 2.34 5.52 6.74 4.79 5.95 8.07
MSE 5.03 4.00 4.42 3.62 3.58 4.38
Data 4 W 1.17 1.99 1.72 4.54 1.98 5.70
MSE 3.26 3.04 3.65 3.19 3.36 3.90
Data 5 W 3.97 3.66 3.03 7.88 7.65 2.54
MSE 4.85 3.56 3.65 6.08 4.01 3.66
Data 6 W 6.15 2.52 7.34 5.98 2.42 7.45
MSE 5.65 4.4 5.62 5.87 4.54 6.44
Data 7 W 5.66 5.10 3.29 2.38 0.99 1.99
MSE 5.20 3.76 3.88 3.71 3.55 3.72
Data 8 W 2.64 0.88 2.32 3.88 3.96 2.23
MSE 3.11 2.78 3.48 3.90 4.12 2.81
Data 9 W 9.28 1.88 3.10 8.74 8.71 4.42
MSE 4.08 3.86 7.43 7.65 5.02 3.71
Data 10 W 1.25 5.40 2.12 0.99 0.57 0.90
MSE 2.88 3.68 2.58 2.64 2.65 2.62
#win W 0 7 6 5 5 6
MSE 0 8 6 5 6 7
Table 2: Calibrating with the multivariate time series ($F_{2}$-$F_{6}$) robustly outperforms that with the univariate data ($F_{1}$).
@@
Calibration of the synthetic data. For each of the 10 synthetic data, by optimizing the 6 objective functions $F_{i},i=1,...,6$ with PSO, it results in 6 calibrated models and 6 simulated data. Note that, even though a model is calibrated using $F_{i},i<6$, it can still functionally generate the simulated data $\mathbf{X}_{T}$ with respect to all 6 features.
Figure 5: The simulated data using the proposed method resembles the synthetic data on all 6 features.
All 6 discrepancy values $D$ between each simulated data and its corresponding synthetic data is measured by both the W and MSE, based on the 6 features, respectively. In Table  2, each cell displays the average of these 6 discrepancy values. It shows that by using more than one feature ($F_{2}$ to $F_{6}$) for calibration, it consistently outperforms the calibration with the univariate data ($F_{1}$). Specifically, on the synthetic data of data 1, 3, and 7, calibrating with $F_{1}$ is defeated by all of the other 5 objective functions who use at least two features. On other instances, there are at least two objectives among $F_{2}$ to $F_{6}$ outperform $F_{1}$.
To a further step, the numbers of synthetic data that each $F_{i},i=2,...,6$ outperforms $F_{1}$ are counted and recorded in the last row of Table  2. It clearly shows that each $F_{i},i=2,...,6$ wins $F_{1}$ on at least half of the testing instances on both W and MSE metrics. More specifically, $F_{2}$ wins the most among $F_{2}$ to $F_{6}$. Note that, all these 5 objectives employ the second feature of total traded volume. It can be attributed that this feature contributes the most to the calibration of multiple features. This is reasonable as the traded volume is a data feature very different to the first feature of mid-price.
@@
Then, we depict the high-frequency time series of 6 features to show how the simulated data resembles the synthetic data in Figure  5. Specifically, the simulated data calibrated by $F_{5}$ on data 7 is displayed as it represents the best calibration performance that $F_{2}$ to $F_{6}$ all outperform $F_{1}$ on both the Wasserstein distance and MSE metric. It can be seen that the simulated time series resemble the synthetic data on all 6 features very closely. To our best knowledge, this is the first ABM-agnostic method of successfully calibrating the high-frequency market time series at 1 second level and simulating 6 features with high-fidelity. In comparison, existing works mostly calibrate the mid-price data at a daily frequency [6].
Calibration of the real data. At last, it is verified that the proposed method can also calibrate the high-frequency real market data at a 3-second level. A 1-hour time series of 000001.sz is calibrated using $F_{1}$, $F_{2}$, and $F_{3}$. The simulated data generated by the corresponding best calibrated models is depicted in Figure  6, together with the real observed data. Due to the limited space, only the mid-price time series is shown in a time step of 1 minute. It can be seen that the simulated mid-price of $F_{1}$ is less similar to the real observed data. By incorporating more features, the fidelity of the simulated mid-price improves by $F_{2}$ and $F_{3}$. This confirms that the proposed method indeed can calibrate the real market data with multiple time series, thus suggesting a new effective calibration objective function for financial market simulation.

### v3

@@ 3.2 Alleviating the Non-identifiability with Multiple Features
Note that, given any two parameters $\omega_{1},\omega_{2}\in\Omega^{D,k}$, only they are simultaneously non-identifiable in all $K$ individual calibration tasks, i.e., $\omega_{1},\omega_{2}\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$, are they non-identifiable in the general calibration task with respect to multivariate $\hat{\mathbf{X}}_{T}$. Otherwise, their discrepancy to $\hat{\mathbf{X}}_{T}$ can be distinguished in at least one feature and thus are identifiable. The search space $\Omega^{D,k}$ is identical across all calibration tasks and is uniformly represented as $\Omega^{D}$. That is to say, only the intersection of the $K$ individual non-identifiable set $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$ can be regarded as the non-identifiable set in the multivariate time series setting, which is defined as
Definition 3 (The multivariate non-identifiability)
Let $P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)$ be the non-identifiability to the multivariate time series with $K$ features. It it defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D}$ while it falls into all $K$ individual non-identifiable set $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$. And we have
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)=\frac{\left|\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D}\right|}.$ (4)
Since $\frac{\left|\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D}\right|}\leq\min_{k}\frac{\left|\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D,k}\right|}$, we have $P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)\leq\min_{k}P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right)$. This equality holds only if the $K$ non-identifiable sets are fully overlapped. This can be caused that the $K$ features are fully dependent that the time series data of one feature can be deduced from other features. This suggests that the selected features should be as diverse as possible. Therefore, with proper choice of the features and $D$, the non-identifiability of calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$ is smaller than that of the univariate $\hat{\mathbf{X}}_{T}^{k}$. In other words, calibrating multiple features can alleviate the non-identifiability issue. Next, we estimate how fast it decreases with $K$ as follows.
For clarity, let $A_{k}$ denote the event of a uniformly randomly sampled parameter falling into $\mathcal{S}^{D,k,\epsilon}$. The probability of happening $A_{k}$ is $P(A_{k})=P(\omega\in\mathcal{S}^{D,k,\epsilon}|\hat{\mathbf{X}}_{T}^{k})$. Then we have:
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)$
$=P\left(A_{1}A_{2}\cdots A_{K}\right)$
$=P(A_{1})\cdot P(A_{2}|A_{1})\cdots P(A_{K}|A_{K-1}\cdots A_{1})$
$=\frac{\left|\mathcal{S}^{D,1,\epsilon}\right|}{\left|\Omega^{D}\right|}\cdot\frac{\left|\cap_{i=1}^{2}\mathcal{S}^{D,i,\epsilon}\right|}{\left|\mathcal{S}^{D,1,\epsilon}\right|}\cdots\frac{\left|\cap_{i=1}^{K}\mathcal{S}^{D,i,\epsilon}\right|}{\left|\cap_{i=1}^{K-1}\mathcal{S}^{D,i,\epsilon}\right|}$
@@
[…]
Theorem 1 (The exponential alleviation)
If each $K$-th feature is selected such that $\beta_{K}^{2}<\frac{{\sum}_{i=1}^{K-1}\beta_{i}^{2}}{K-1}$, the upper bound of the multivariate non-identifiability is reduced exponentially with the increasing number of features $K$.
[…]
@@ 3.3 Aggregating Multiple Features via Maximization
Note that, the above discussions require the multivariate time series $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$ being calibrated separately as $K$ univariate time series with existing discrepancy metrics $D$. This results in $K$ individual calibration tasks of $D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),k=1,\cdots,K$. How to jointly calibrate the parameter to satisfy the $K$ individual tasks simultaneously? This section proposes a new objective function based on the $K$ individual objectives, so that the practical alleviation of the non-identifiability follows the above theory.
Let $F$ be the proposed new objective function used for calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$, and let $\mathcal{S}^{F,\epsilon}$ denote the associated non-identifiable set defined by this objective function. To achieve $\mathcal{S}^{F,\epsilon}=\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$, the proposed way is to ensure every location in $S^{F}$ to be the worst value among all $K$ non-identifiable sets $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$ (see Figure 3 for illustration). In this case, any parameter $\omega$ that is not simultaneously in all $K$ non-identifiable sets will make it fall outside of the $\mathcal{S}^{F,\epsilon}$. This is because that the parameters outside the non-identifiable set is even worse than those who are non-identifiable from the ground-truth parameter $\omega^{*}$.
Figure 3: Optimizing the maximization of two functions equals to searching in their intersections.
Given that the calibration problem is a minimization problem, the new objective function $F$ can be implemented with a maximization aggregation among $K$ individual calibration tasks with $D$. Formally, we have the following theorem.
Theorem 2 (The utility of new objective function $F$)
Let $F$ aggregate the $K$ individual calibration tasks via maximization, i.e.
$F(\hat{\mathbf{X}}_{T},M(w)=\mathbf{X}_{T})$ (6)
$=$ $\max_{k\in\{1,\cdots,K\}}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),$
then minimizing $F$ achieves $\mathcal{S}^{F,\epsilon}=\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$.
The proof of Theorem 2 is provided in appendix.
@@ 4 Empirical Studies
This section empirically verifies two major claims.
  • With the increasing number of features applied in the calibration, the non-identifiable set shrinks exponentially.
@@ 4.1 Experimental Settings
$D(\hat{\mathbf{X}}_{T}^{k},\mathbf{X}_{T}^{k})$ (7)
$=$ $\int_{-\infty}^{+\infty}\left|\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\hat{\mathbf{X}}_{t}^{k}\leq x)-\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\mathbf{X}_{t}^{k}\leq x)\right|\,\mathrm{d}x,$
where $\mathcal{I}(\cdot)$ is an indicator function that returns 1 if the input event is true, otherwise it returns 0.
The simulation model. This work considers the Preis-Golke-Paul-Schneid (PGPS) model as the simulation model, due to its popularity in recent financial market simulation works  [25, 11]. The PGPS model simulates the market dynamics through interactions between two types of agents, i.e., 125 liquidity providers and 125 liquidity takers. The PGPS model has 6 hyper-parameters shared by all agents that need to be calibrated, i.e., $\omega=[\delta,\lambda_{0},C_{\lambda},\Delta_{s},\alpha,\mu]$. The detailed descriptions of the behaviors of the agents and how the hyper-parameters control them can be found in Appendix.
The optimization method. Since the order matchmaking process needs to align with the real market, the PGPS model should be run on the executable trading programs. As a result, the calibration objective function is non-differentiable and the calibration process often has to resort to non-derivative optimization methods. In such contexts, both Bayesian optimization methods and evolutionary algorithms are particularly suitable. This work employs the well-known Particle Swarm Optimization (PSO) algorithm  [15] to optimize $\omega$ with respect to Eq. (6). The setting of PSO is in Appendix.
@@
  • $f_{3}$, the price return at each $t$-th step $\ln m_{t+1}-\ln m_{t}$;
@@
The test protocol. The test data contains 11 targeted time series to be calibrated, including 10 synthetic data and 1 real data. The 10 synthetic time series are generated by setting 10 different parameters of the PGPS model as shown in Appendix, each of which consists of 3600 time steps at the frequency of 1 second. The real data is chosen as 000001.sz from Shenzhen Stock Exchange of China, consisting of 1200 time steps at a frequency of 3 seconds from 9:30 a.m. to 10:30 a.m. of a day in 2019. The time budget of $10^{4}$ simulation evaluations is allowed for the calibration of each targeted data. That is, the PSO runs for 250 iterations with population size 40, and the best parameter with the lowest objective function value is applied into PGPS model to simulate data as $\mathbf{X}_{T}$. The performance of the calibration is quantitatively validated using the Wasserstein (W) distance and the Mean Square Error (MSE) commonly seen in time series analysis. Both metrics are applied to measure the averaged discrepancy between the synthetic data and simulated data on 6 features, i.e., $\frac{\sum_{k=1}^{6}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k})}{6}$, where $D$ here indicates W or MSE. Each feature is normalized with the min-max method, respectively.
@@ 4.2 Results and Analysis
Figure 4: The illustration of the exponentially decreased non-identifiable set (grey dots). The red dot is the optima.
@@
Alleviation of non-identifiability. Figure 4 illustrates the changes of the non-identifiable sets in the 2-dimensional parameter space of $[\alpha,\mu]$. This is done by three steps: 1) a target data of 3600 time steps is generated using the PGPS model with the recommended default parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$  [11]; 2) On this default setting, 10000 combinations are created by sampling in $[\alpha,\mu]$ in grid while keeping the values of other 4 parameters unchanged; 3) By using $F$ to measure the descrepancy between the target data and the simulated data of each of the 10000 combinations, it obtains the objective values of the 10000 combinations on each $F_{i}$, which are depicted as Figure 4. The 6 figures in Figure 4 are depicted according to the objective values of $F_{1}$, $F_{2}$,…,$F_{6}$, respectively. Only the combinations whose objective values smaller than $\epsilon=0.1$ are depicted in grey, otherwise they will be depicted in blue. For example, in the third figure marked with $F_{3}$, the combinations depicted in grey are those whose discrepancy values smaller than $\epsilon=0.1$ on all the first 3 features $\{f_{i}\}_{i=1}^{3}$. By this means, the grey points indicate the non-identifiable set as they are indistinguishable with $\omega^{*}$ (marked as the red dot). As can be seen that, with the increase of the features, the grey non-identifiable set, decreases significantly. When 4 are incorporated in $F$, there exists only the optimum, indicating that the non-identifiability issue is indeed alleviated very effectively.
[…]
Figure 5: The non-identifiability (NI) decreases exponentially with the increasing features.
Figure 6: The distances of the simulated data to the target data decrease almost linearly with the increasing number of features incorporated in $F$.
@@
Calibration to the synthetic data. Note that, although the non-identifiable set of any $F_{i}$ is verified to be smaller than that of any $F_{j}$, $i>j$, it does not deterministically lead to better calibration performance of $F_{i}$ than $F_{j}$. Because $F$ mainly decides the parameter space, while the calibration performance is also jointly influenced by the optimization algorithms. Given the calibration problem is black-box with multiple local optima, it is non-trivial for the optimization algorithms to find the global optimum even the parameter space is modeled fully identifiable.
@@
For each of the 10 synthetic data, by optimizing the 6 objective functions $F_{i},i=1,...,6$ with PSO, it results in 6 calibrated models and 6 simulated data. Note that, even though a model is calibrated using $F_{i},i<6$, it can still functionally generate the simulated data $\mathbf{X}_{T}$ with respect to all 6 features.
The performance between each simulated data and its corresponding synthetic data is measured by the indicators based on both MSE and W, respectively. In Figure 6, each (blue or yellow) point displays the average performance on these 10 synthetic data. It shows that by using more features in $F$, the performance of calibration generally improves as the distance values decrease.
Furthermore, it is interesting to see how the simulation model calibrated by all features ($F_{6}$) compares to the ones calibrated to each single feature $f_{i},i=1,...,6$, instead of in an incremental manner of $F_{i},i=1,...,5$. As listed in Table 1, by calibrating with $F_{6}$, the obtained parameter enjoys the best simulation performance, i.e., the smallest averaged distances on 7 out of the 10 instances, and top 3 averaged distances on the other 3 instances. Besides, the friedman test is applied to show that $F_{6}$ achieves the best average performance and the advantage is statistically significant. This implies that the proposed theory and new objective function are applicable to various features and not restricted by the feature with more information of the social system, e.g., the mid-price.
$f_{1}$ $f_{2}$ $f_{3}$ $f_{4}$ $f_{5}$ $f_{6}$ $F_{6}$
Data 1 3.99 5.07 6.22 4.62 4.92 3.95 3.07
Data 2 3.65 3.95 4.67 3.73 6.18 4.77 2.90
Data 3 2.86 4.43 3.73 3.87 5.75 4.62 3.08
Data 4 2.83 3.50 3.47 3.87 4.33 2.99 2.47
Data 5 2.72 2.70 4.93 5.98 3.17 5.83 2.91
Data 6 5.23 5.58 4.02 6.67 4.48 6.77 3.67
Data 7 3.08 3.57 3.06 3.65 3.60 3.86 3.51
Data 8 4.88 4.50 2.56 3.82 3.91 3.62 2.26
Data 9 4.31 3.59 5.07 4.19 5.73 3.82 3.51
Data 10 2.53 2.43 2.95 6.93 3.22 2.24 2.23
#rank 3.20 3.60 3.80 4.50 5.30 3.80 1.57
Table 1: Comparisons of calibration performance between 6 univariate data and $F_{6}$. Note that the performance is at the $1E10^{-2}$ level.
Then, we illustrate the time series of 6 features to show how the simulated data resembles the synthetic data in Figure 7. Specifically, we calibrate the model by $F_{6}$ and $f_{x}$ on the data generated based on the recommended parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$  [11], where $f_{x}$ indicates any single feature who helps to calibrate the model and generate the best simulated data. And the simulated data calibrated by $F_{6}$ resembles the target data on all 6 features very closely. To our best knowledge, this is the first ABM-agnostic method of successfully calibrating the high-frequency market time series at 1 second level and simulating 6 features with high-fidelity. In comparison, the simulated data of $f_{x}$ performs much poorer.
Figure 7: The simulated data using $F_{6}$ resembles the synthetic data on all 6 features.
Calibration to the real data. At last, it is verified that the proposed method can also calibrate the real market data at a 3-second level. A 1-hour time series of 000001.sz is calibrated using $F_{1}$ and $F_{3}$. The simulated data generated by the corresponding best calibrated models is depicted in Figure 8, together with the real observed data. Note that, most of the FMS simulators (including the adopted PGPS model) model the order volume in a coarse-grained manner, i,e., simply setting all the orders with the identical volume of 100. Hence, only the mid-price time series is shown in a time step of 1 minute. It can be seen that the simulated mid-price of $F_{3}$ is more similar to the real observed data than $F_{1}$ by an improvement of 27.5% in MSE. This confirms that the proposed method indeed can calibrate the real market data with multiple time series, thus suggesting a new effective calibration objective function for financial market simulation.